# compat=1.0.0

### v1

@@ III Thermal production of charm quarks in the QGP
Now we compute the thermal production of charm pairs in heavy-ion collisions at RHIC and LHC energies. For this purpose the PHSD is used to simulate heavy-ion collisions Song et al. (2015, 2016) where charm is produced only through initial hard scattering of nucleons. The energy-momentum of (anti)charm quarks is provided by the PYTHIA event generator Sjostrand et al. (2006), in which the transverse momentum and rapidity are tuned such that they become similar to those in the Fixed-Order Next-to-Leading Logarithm (FONLL) calculations Cacciari et al. (2012). In heavy-ion collisions, the charm energy-momentum distribution is also modified due to the (anti)shadowing effects which are realized in PHSD by EPS09 Eskola et al. (2009). We recall that charm production is suppressed at low transverse momentum and at mid-rapidity which arise from small $x$ of the parton distribution function Song et al. (2016).
@@
Since neither thermal production nor annihilation of charm quarks is considered, the number of charm is conserved in heavy-ion collisions. However, including the processes in Figs. 1 and 2 for charm production as well as charm annihilation in the QGP, the charm number is not conserved any more.
@@
Figure 4: (Top) The PHSD results for the numbers of charm pair production and annihilation per rapidity bin at mid-rapidity as a function of time (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. The charm quark mass is taken to be 1.5 GeV.
@@
Figure 5: The distributions of scattering energy $\sqrt{s}$ for thermal charm production per rapidity bin at mid-rapidity for PHSD and for the linearized Boltzmann approach in the PHSD assuming local thermal equilibrium (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. The upper and lower limits of the bands indicate the thermal charm production from $t=$ 0 $fm/c$ and 0.5 $fm/c$, respectively, in the linearized Boltzmann approach. The charm quark mass is taken to be 1.5 GeV.
@@
Fig. 4 shows the number of thermally produced and annihilated charm quark pairs per rapidity bin at midrapidity as a function of time in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV for a charm quark mass of 1.5 GeV. One can see that the thermal production of charm dominantly takes place in the early stage of heavy-ion collisions where the temperature is high and partons have larger energy to produce a charm quark pair. If the partonic matter is completely thermalized, the contributions from quark-antiquark annihilation and from two gluon fusion will be similar. But the initial stage of heavy-ion collisions is far off equilibrium. In the literature different types for off-equilibrium are quoted Song et al. (2020a), for example, the chemical off-equilibrium, the kinetic off-equilibrium, the anisotropic off-equilibrium and the off-equilibrium in parton spectral functions. The PHSD accounts for all of them.
@@
In the PHSD initial partons are produced by string melting when the local energy density exceeds a critical value which is $\sim 0.5~{}{\rm GeV/fm^{3}}$. The mass distributions of quarks and gluons follow the spectral functions from the DQPM through scatterings. We mention that the initial mass distributions of quarks and gluons are different from the thermalized spectral distributions of the DQPM. In PHSD gluons are initially more massive, but (anti)quarks are less massive compared to the equilibrium spectral functions of the DQPM. Thus, the thermal production of charm quark pairs is dominated by two gluon fusion in the early stage of heavy-ion collisions as shown in Fig. 4. We also note that initial partons, which are produced by the string melting in PHSD, have a formation time, which is proportional to $E/m_{T}^{2}$ with $E$ and $m_{T}$ being, respectively, energy and transverse mass of the parton; the parton does not interact before the formation time has passed. That is why the charm production is suppressed at the very early time ($\leq 0.2$ fm/c) in Fig. 4.
@@
When the local temperature decreases and approaches $T_{c}$, all gluons split into quark and antiquark pairs which is followed by quark coalescence to form mesons and (anti)baryons. Therefore two gluon fusion is suppressed and quark-antiquark annihilation becomes the dominant channel for thermal charm quark production, although the production rate is much smaller than in the initial gluon-gluon fusion. The number of produced thermal charm pairs per rapidity at mid-rapidity is about 0.15 in central Au+Au collisions at 200 GeV and a 2.1 in central Pb+Pb collisions at 5.02 TeV.
@@
On the other hand charm quark annihilation, which is realized by detailed balance, contributes similarly to the quark and antiquark production and two gluon production as shown by the dotted lines in Fig. 4, but they are less than the charm production channels.
@@
Fig. 5 shows the scattering energy distribution which produces thermal charm quark pairs at mid-rapidity in central heavy-ion collisions at RHIC and LHC energies. In order to see the non-equilibrium effects of partonic matter on charm production, the results from PHSD are compared with those from the linearized Boltzmann approach Song et al. (2020b), shown as colored bands for quark-antiquark annihilation and two gluon fusion channels. The upper limit of the band indicates thermal charm production from $t=$ 0 fm/$c$ assuming thermal equilibrium as soon as the two nuclei pass through each other, and the lower limit is for $t=$ 0.5 fm/$c$, which is a typical initial thermalization time for hydrodynamical simulations at RHIC and LHC energies. One can see that the results from the linearized Boltzmann approach are very sensitive to the initial thermalization time, because the temperature in the initial stage is high and more effective to produce thermal charm quarks. The colored bands are wider at RHIC, because the lifetime of the QGP is relatively short and the initial time period between $t=$ 0 and 0.5 fm/$c$ is more important than at LHC. The broad bands in the figure show the difficulty of studying thermal charm production in a hydrodynamical approach since in this calculations charm production before the initial thermalization time is not included.
@@
Figure 6: (Upper) The PHSD results for the numbers of charm quark pairs produced initially and thermally at mid-rapidity as a function of time and (lower) their $p_{T}$ spectra in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV from the PHSD. The charm quark mass is taken to be 1.5 GeV.
@@
In the upper panel of Fig. 6 we show the number of initially and thermally produced charm quark pairs at midrapidity as a function of time in central Au+Au and Pb+Pb collisions at RHIC and LHC. The initial production caused by primary nucleon-nucleon binary scattering takes place shortly after the two nuclei overlap. On the other hand, the thermal production takes time and its number increases gradually, although most of them are produced before $t\simeq 1-2$ fm/$c$. One can see that the number of thermal charm quarks at LHC is comparable to that of initial charm quarks at RHIC for $m_{c}$= 1.5 $\rm{GeV}$.
@@
The lower panel shows the $p_{T}$ spectra of initially produced charm quark pairs and those of thermal charm quark pairs for $m_{c}=$ 1.5 $\rm{GeV}$ at mid-rapidity in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. We note that these are spectra at the production time and the interactions in the QGP are not included. Comparing the solid and dashed lines, the spectra of the thermal charm are softer than those of the initial hard charm pairs. Counting the numbers, the number of thermal charm in mid-rapidity is about 6.5 % of the number of initially produced charm at the same rapidity at RHIC and 18 % at LHC.
@@
Figure 7: The PHSD results for the $R_{\rm AA}$ of $D$ mesons at mid-rapidity as a function of transverse momentum (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV with and without thermal charm production in comparison to the experimental data from the STAR Adamczyk et al. (2014); Adam et al. (2019) and ALICE Acharya et al. (2022) Collaborations. The charm quark mass for thermal production varies from 1.2 GeV to 1.8 GeV which corresponds, respectively, to the upper and lower limits of the colored bands.
@@
Finally, we examine the effects of thermal charm - including the interactions of charm quarks in the QGP - on experimental observables in heavy-ion collisions such as the nuclear modification of $D$ mesons $\rm R_{AA}$ which is defined as
${\rm R_{AA}}(p_{T})\equiv\frac{dN_{D}^{\rm A+A}/dp_{T}}{N_{\rm binary}^{\rm A+A}\times dN_{D}^{\rm p+p}/dp_{T}},$
@@
where $N_{D}^{\rm A+A}$ and $N_{D}^{\rm p+p}$ are, respectively, the number of $D$ mesons produced in heavy-ion collisions and that in p+p collisions, and $N_{\rm binary}^{\rm A+A}$ is the number of binary nucleon-nucleon collisions in heavy-ion collision for the centrality class considered.
@@
Figure 7 shows the $R_{\rm AA}$ of $D$ mesons including thermal charm production at mid-rapidity as a function of transverse momentum (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV compared to the experimental data from the STAR Adamczyk et al. (2014); Adam et al. (2019) and ALICE Acharya et al. (2022) Collaborations. So far the charm quark mass has been taken to be 1.5 GeV, however, the thermal production of charm pairs depend on the charm quark mass since the threshold is $2m_{c}$. If charm quark is lighter, its thermal production will increase. We vary the charm quark mass from 1.2 GeV, which is close to the bare mass in QCD, to 1.8 GeV, which is close to the $D$ meson mass. The upper limit of the bands in Fig. 7 corresponds to the $R_{\rm AA}$ of $D$ meson with the thermal charm quark mass being 1.2 GeV and the lower limits for 1.8 GeV. As shown in Fig. 7 the thermal charm quarks enhance $R_{\rm AA}$ at small transverse momentum and have only small effect at large transverse momentum. We note that the discrepancies at large $p_{T}$ between the solid line and colored band are not the effects of thermal charm quarks but merely statistical fluctuations. The number of thermal charm quark pairs at mid-rapidity is 0.7--0.053 per rapidity at RHIC and 11.5--0.55 at LHC for $m_{c}=1.2-1.8$ GeV.
@@
Based on Fig. 7 it is hard to say something robust about the thermal charm quarks effects $R_{\rm AA}$ on at RHIC because the contribution from the thermal charm is small compared to the uncertainties in the experimental data. On the other hand, the effects of thermal charm quarks are more visible at LHC because the error bars of experimental data are relatively small and more thermal charm pairs are produced. The $R_{\rm AA}$ of $D$ mesons including thermal charm for $m_{c}=1.8$ GeV is almost the same as the results without the thermal production. For $m_{c}=1.2$ GeV, however, the effects are clearly visible and the $R_{\rm AA}$ of $D$ mesons is enhanced at low transverse momentum. Compared with the experimental data from the ALICE Collaboration, the small charm quark mass of 1.2 GeV overestimates the experimental data. This suggests that a larger charm quark mass is more probable for the thermal production in heavy-ion collisions, which is consistent with the gaining mass in the QGP through particle dressing Moreau et al. (2019); Gubler et al. (2020).
@@
Figure 8: The DQPM results for the scaled spatial diffusion coefficient of charm quarks as a function of temperature ($T_{c}=$ 0.158 GeV) for charm quark masses ranging from 1.2 GeV to 1.8 GeV, which correspond to the upper and lower limits of the orange band. The coefficients are compared with those from lattice calculations Banerjee et al. (2012); Altenkort et al. (2024).
@@
Finally we discuss the effect of a charm quark mass on its spatial diffusion coefficient in the QGP. Fig. 8 shows the spatial diffusion coefficient of charm quarks as a function of temperature $T$ scaled by the critical temperature $T_{c}$. The upper and lower limits of the orange band correspond to $m_{c}=$ 1.2 GeV and 1.8 GeV, respectively. The diffusion coefficient decreases with increasing charm quark mass. The orange band is compared with two lattice calculations Banerjee et al. (2012); Altenkort et al. (2024) one of which is consistent with our results, while the other is about a factor of 2 down. A smaller diffusion coefficient can be obtained by increasing the strong coupling for charm quark elastic scattering and it will enhance thermal production of charm quarks. Therefore, it strongly supports our conclusion that the experimental data favor heavier charm quark than its bare mass.

### v4

@@ IV Thermal production of charm quark pairs in heavy-ion collisions
Now, we compute the thermal production of charm pairs in heavy-ion collisions at RHIC and LHC energies. For this purpose, the PHSD is used to simulate heavy-ion collisions Song et al. (2015, 2016), where charm is produced only through initial hard scattering of nucleons. The energy-momentum of (anti)charm quarks is provided by the PYTHIA event generator Sjostrand et al. (2006), in which the transverse momentum and rapidity are tuned to match those in the Fixed-Order Next-to-Leading Logarithm (FONLL) calculations Cacciari et al. (2012). In heavy-ion collisions, the charm energy-momentum distribution is also modified due to the (anti)shadowing effects, which are realized in PHSD by EPS09 Eskola et al. (2009). We recall that charm production is suppressed at low transverse momentum and at midrapidity, which arises from small $x$ of the parton distribution function Song et al. (2016).
@@
Since neither thermal production nor annihilation of charm quarks is considered, the number of charm quarks is conserved in heavy-ion collisions. However, by including the processes in Figs. 2 and 3 for charm production as well as charm annihilation in the QGP, the charm number is not conserved anymore.
@@
Figure 5: The PHSD results for the numbers of charm pair production and annihilation per rapidity bin at midrapidity as a function of time (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. The charm quark mass is taken to be 1.5 GeV.
@@
Figure 6: The distributions of scattering energy $\sqrt{s}$ for thermal charm production per rapidity bin at midrapidity for PHSD and for the linearized Boltzmann approach in the PHSD assuming local thermal equilibrium (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. The upper and lower limits of the bands indicate the thermal charm production from $t=$ 0 $fm/c$ and 0.5 $fm/c$, respectively, in the linearized Boltzmann approach. The charm quark mass is taken to be 1.5 GeV.
@@
Figure 5 shows the number of thermally produced and annihilated charm quark pairs per rapidity bin at midrapidity as a function of time in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV for a charm quark mass of 1.5 GeV. One can see that the thermal production of charm dominantly takes place in the early stage of heavy-ion collisions, where the temperature is high and partons have larger energy to produce a charm quark pair. If the partonic matter is completely thermalized, the contributions from quark-antiquark annihilation and from two-gluon fusion will be similar. However, the initial stage of heavy-ion collisions is far from equilibrium. In the literature, different types of off-equilibrium are quoted Song et al. (2020a), for example, the chemical off-equilibrium, the kinetic off-equilibrium, the anisotropic off-equilibrium, and the off-equilibrium in parton spectral functions. The PHSD accounts for all of them.
@@
In the PHSD, initial partons are produced by string melting when the local energy density exceeds a critical value of approximately 0.5 GeV/fm3. The mass distributions of quarks and gluons follow the spectral functions from the DQPM through scatterings. We note that the initial mass distributions of quarks and gluons differ from the thermalized spectral distributions of the DQPM. In the PHSD, gluons are initially more massive, but (anti)quarks are less massive compared to the equilibrium spectral functions of the DQPM. Thus, the thermal production of charm quark pairs is dominated by two-gluon fusion in the early stage of heavy-ion collisions, as shown in Fig. 5. We also note that initial partons, which are produced by string melting in the PHSD, have a formation time that is proportional to $E/m_{T}^{2}$, with $E$ and $m_{T}$ being energy and transverse mass of the parton, respectively; the parton does not interact before the formation time has passed. This is why the charm production is suppressed at a very early time ($\leq 0.2$ fm/c) in Fig. 5.
@@
When the local temperature decreases and approaches $T_{c}$, all gluons split into quark and antiquark pairs, followed by quark coalescence to form mesons and (anti)baryons. Therefore, two-gluon fusion is suppressed, and quark-antiquark annihilation becomes the dominant channel for thermal charm quark production, although the production rate is much smaller than in the initial gluon-gluon fusion. The number of produced thermal charm pairs per unit rapidity at midrapidity is about 0.15 in central Au+Au collisions at 200 GeV and 2.1 in central Pb+Pb collisions at 5.02 TeV.
@@
On the other hand, charm quark annihilation, which is realized by detailed balance, contributes similarly to the quark-antiquark production and two-gluon production as shown by the dotted lines in Fig. 5, but these contributions are less than those of the charm production channels.
@@
Figure 6 shows the scattering energy distribution, which produces thermal charm quark pairs at midrapidity in central heavy-ion collisions at RHIC and LHC energies. To observe the non-equilibrium effects of partonic matter on charm production, the results from PHSD are compared with those from the linearized Boltzmann approach Song et al. (2020b), shown as colored bands for quark-antiquark annihilation and two-gluon fusion channels. The upper limit of the band indicates thermal charm production from $t=$ 0 fm/$c$, assuming thermal equilibrium as soon as the two nuclei pass through each other, and the lower limit is for $t=$ 0.5 fm/$c$, which is a typical initial thermalization time for hydrodynamical simulations at RHIC and LHC energies. One can see that the results from the linearized Boltzmann approach are highly sensitive to the initial thermalization time because the temperature in the initial stage is high and more effective to produce thermal charm quarks. The colored bands are wider at RHIC because the lifetime of the QGP is relatively short, and the initial time period between $t=$ 0 and 0.5 fm/$c$ is more important than at the LHC. The broad bands in the figure show the difficulty of studying thermal charm production in a hydrodynamical approach since in these calculations charm production before the initial thermalization time is not included.
@@
Figure 7: (Upper) The PHSD results for the numbers of charm quark pairs produced initially and thermally at midrapidity as a function of time and (lower) their $p_{T}$ spectra in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV from the PHSD. The charm quark mass is taken to be 1.5 GeV.
@@
In the upper panel of Fig. 7, we show the number of initially and thermally produced charm quark pairs at midrapidity as a function of time in central Au+Au and Pb+Pb collisions at RHIC and the LHC. The initial production, caused by primary nucleon-nucleon binary scattering, takes place shortly after the two nuclei overlap. On the other hand, thermal production takes time, and its number increases gradually, although most are produced before $t\simeq 1-2$ fm/$c$. One can see that the number of thermal charm quarks at the LHC is comparable to that of initial charm quarks at RHIC for $m_{c}$= 1.5 GeV.
@@
The lower panel shows the $p_{T}$ spectra of initially produced charm quark pairs and those of thermal charm quark pairs for $m_{c}=$ 1.5 GeV at midrapidity in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. We note that these are spectra at the production time and that the interactions in the QGP are not included. Comparing the solid and dashed lines, the spectra of the thermal charm are softer than those of the initial hard charm pairs. Counting the numbers, the number of thermal charm quarks at midrapidity is about 6.5% of the number of initially produced charm quarks at the same rapidity at RHIC and 18% at the LHC.
@@
Figure 8: The PHSD results for the $R_{\rm AA}$ of $D$ mesons at midrapidity as a function of transverse momentum (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV with and without thermal charm production in comparison to the experimental data from the STAR Adamczyk et al. (2014); Adam et al. (2019) and ALICE Acharya et al. (2022) Collaborations. The charm quark mass for thermal production varies from 1.2 GeV to 1.8 GeV, which corresponds, respectively, to the upper and lower limits of the colored bands.
@@
Finally, we examine the effects of thermal charm – including the interactions of charm quarks in the QGP – on experimental observables in heavy-ion collisions such as the nuclear modification of $D$ mesons $\rm R_{AA}$, which is defined as
$R_{\mathrm{AA}}(p_{T})\equiv\frac{dN_{D}^{\rm A+A}/dp_{T}}{N_{\rm binary}^{\rm A+A}\times dN_{D}^{\rm p+p}/dp_{T}},$
@@
where $N_{D}^{\rm A+A}$ and $N_{D}^{\rm p+p}$ are, respectively, the number of $D$ mesons produced in heavy-ion collisions and in p+p collisions, and $N_{\rm binary}^{\rm A+A}$ is the number of binary nucleon-nucleon collisions in heavy-ion collision for the considered centrality class.
@@
Figure 8 shows the $R_{\rm AA}$ of $D$ mesons including thermal charm production at midrapidity as a function of transverse momentum (upper) in central Au+Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV and (lower) in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV, compared to the experimental data from the STAR Adamczyk et al. (2014); Adam et al. (2019) and ALICE Acharya et al. (2022) collaborations. So far, the charm quark mass has been taken to be 1.5 GeV. However, the thermal production of charm pairs depends on the charm quark mass due to the threshold of $2m_{c}$. If the charm quark is lighter, its thermal production will increase. We vary the charm quark mass from 1.2 GeV, which is close to the bare mass in QCD, 1.8 GeV, which is close to the $D$ meson mass. The upper limit of the bands in Fig. 8 corresponds to the $R_{\rm AA}$ of the $D$ meson with the thermal charm quark mass being 1.2 GeV, and the lower limits to 1.8 GeV. As shown in Fig. 8, the thermal charm quarks enhance $R_{\rm AA}$ at small transverse momentum and have only a small effect at large transverse momentum. We note that the discrepancies at large $p_{T}$ between the solid line and colored band are not the effects of thermal charm quarks but merely statistical fluctuations. The number of thermal charm quark pairs at midrapidity is 0.7--0.053 per unit rapidity at RHIC and $11.5-0.55$ at the LHC for $m_{c}=1.2-1.8$ GeV.
@@
Based on Fig. 8, it is hard to say something robust about the thermal charm quark effects on $R_{\rm AA}$ at RHIC, as the contribution from the thermal charm is small compared to the uncertainties in the experimental data. On the other hand, the effects of thermal charm quarks are more visible at the LHC because the error bars of the experimental data are relatively small and more thermal charm pairs are produced. The $R_{\rm AA}$ of $D$ mesons including thermal charm for $m_{c}=1.8$ GeV is almost the same as the results without thermal production. For $m_{c}=1.2$ GeV, however, the effects are clearly visible, and the $R_{\rm AA}$ of $D$ mesons is enhanced at low transverse momentum. When compared with the experimental data from the ALICE Collaboration, the small charm quark mass of 1.2 GeV overestimates the experimental data. This suggests that a larger charm quark mass is more probable for thermal production in heavy-ion collisions, which is consistent with the gaining mass in the QGP through particle dressing Moreau et al. (2019); Gubler et al. (2020).
@@
Figure 9: The DQPM results for the scaled spatial diffusion coefficient of charm quarks as a function of the scaled temperature ($T_{c}=$ 0.158 GeV) for charm quark masses ranging from 1.2 GeV to 1.8 GeV, which correspond to the upper and lower limits of the orange band. The coefficients are compared with those from lattice calculations Banerjee et al. (2012); Altenkort et al. (2024).
@@
Finally, we discuss the effect of the charm quark mass on its spatial diffusion coefficient in the QGP. Figure 9 shows the spatial diffusion coefficient of charm quarks as a function of temperature $T$ scaled by the critical temperature $T_{c}$. The upper and lower limits of the orange band correspond to $m_{c}=$ 1.2 GeV and 1.8 GeV, respectively. The diffusion coefficient decreases with increasing charm quark mass. The orange band is compared with two lattice calculations Banerjee et al. (2012); Altenkort et al. (2024), one of which is consistent with our results, while the other is about a factor of two lower. A smaller diffusion coefficient can be obtained by increasing the strong coupling for charm quark elastic scattering, which will enhance the thermal production of charm quarks. Therefore, this strongly supports our conclusion that the experimental data favor a heavier charm quark than its bare mass.
Accounting for next-to-leading order $2\to 3$ processes within the non-perturbative DQPM framefork Grishmanovskii et al. (2024) and a comparison to the pQCD results of Ref. Zhang et al. (2008) would be an interesting subject for the next study. Now we only can speculate that the radiative processes will enhance charm production in heavy-ion collisions and consolidate our finding that the experimental data favor more massive charm quarks in the QGP.